OpenStudy (anonymous):

solve : show work please 2(n^2+n)-5(n^2-4n)

OpenStudy (anonymous):

2n^2+2n-5n^2+20n

OpenStudy (dumbcow):

can't solve without an "="

OpenStudy (anonymous):

-3n^2+22n

OpenStudy (anonymous):

true i think he means simplify looooooool though

OpenStudy (dumbcow):

i know, but its good to know the difference

OpenStudy (anonymous):

just factor it out 2n^2+2n-5n^2+20n then add like terms -3n^2+22n

OpenStudy (anonymous):

6m-5=m+5 6m-m=5+5 5m=10 5m/5=10/5 m=2